OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)=2/x and g(x)=2/x

OpenStudy (whpalmer4):

To show that \(f(g(x)) = x\) you take the definition of \(g(x)\) and substitute it for \(x\) wherever you see \(x\) in the definition of \(f(x)\)

OpenStudy (anonymous):

I say that f and g are inverses because if you take 2 or 4 and plug it in both ways you will get the same answer

OpenStudy (anonymous):

5/4 = x+14/2x-2 If you cross multiply you get 8x -8 = 5x + 70 3x -8 = 70 3x = 78 x = ?

OpenStudy (anonymous):

26

OpenStudy (anonymous):

Ah, thank you both!